Company X has profit margin of 3%, total assets turnover of 1.7, and total debt to total assets ratio of 30%, what is the company's ROE? 
Answer #1

Profit margin is    3.00%  
It means salea is 1, then profit =    0.03  
      
Total Assets turnover=   1.7  
Asset turnover ratio = Total sales/Assets      
1.7 = 1/Total assets      
Total Assets = 1/1.7=   0.5882352941  

Debt is 30%

Equity = 1-30% = 70% of total assets

So Equity =0.5882352941*70%

=0.4117647059

RoE = net income/Equity

0.03/0.4117647059

=0.0729

So RoE is 0.0729

(Please thumbsup)